# TreeFive.bib

### v1

@@ 3. Proof of main result
By Lemma 3.8, there exists a partition $\lambda\vdash n$ with all parts in the interval $J=I=\{b+1,\ldots,b+c\}$. All parts of $\lambda$ are at least $b+1\geq c_{1}+1$ and trivially every rearrangement $\alpha$ of $\lambda$ has the partial sum $\alpha_{1}\in\{b+1,\ldots,b+c\}$, so the result follows from Lemma 3.5. ∎
We can now prove Parts 1 and 2 of Theorem 3.1.
Proposition 3.11.
If $c\geq 2$ and $b\leq 2c-2$, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
@@
If $b\leq 2c-2$, then we have
so the result follows from Proposition 3.10. ∎
Proposition 3.12.
If $c\geq c_{1}+1$ and $b=2c-1$, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
Proof.
First note that if $c=2$, then $b=3$, $I=\{4,5\}$, and taking $\lambda=(2,\ldots,2)$ if $n$ is even or $\lambda=(3,2,\ldots,2)$ if $n$ is odd, every rearrangement $\alpha$ has $\alpha_{1}+\alpha_{2}=4$ or $5$ and we are done by Lemma 3.5. Now suppose that $c\geq 3$. Let $x=\left\lceil\frac{b+1}{2}\right\rceil=c\geq c_{1}+1$ and $y=\left\lfloor\frac{b+c}{2}\right\rfloor=\left\lfloor\frac{3c-1}{2}\right\rfloor$. By Lemma 3.8, we can find a partition $\lambda\vdash n$ with all parts in $J=\{x,\ldots,y\}$ because
By Lemma 3.7 with $q=2$, $G$ has no connected partition of type $\lambda$. ∎
Our last case to prove is Part 3, where $2c\leq b\leq\frac{c^{2}}{2}$. We first check several cases by computer.
Proposition 3.13.
If $2\leq c\leq 40$ and $2c\leq b\leq\frac{c^{2}}{2}$, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
Proof.
By Proposition 3.10, we may assume that $2b+c+1\leq n\leq\left\lceil\frac{b}{c-1}\right\rceil(b+1)$, so there are finitely many triples $(b,c,n)$ to consider. For each one, we checked by computer that there exists a positive integer $q$ and a partition $\lambda\vdash n$ with all parts in the corresponding interval $J=\{x,\ldots,y\}$, where $x\geq c+1\geq c_{1}+1$, so the result follows from Lemma 3.7. ∎
Lemma 3.14.
Suppose that we can find a positive integer $q$ such that the integers $x$ and $y$ as in (3.1) satisfy $x\geq c+1$, $x<y$, and
@@
Then $G$ is missing a connected partition of some type $\lambda\vdash n$.
Proof.
This follows directly from Lemma 3.7 and Lemma 3.8. ∎
Proposition 3.15.
If $41\leq c\leq 500$ and $2c\leq b\leq\frac{c^{2}}{2}$, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
Proof.
There are finitely many pairs $(b,c)$ to consider. For each one, we checked by computer that there exists a positive integer $q$ such that $x\geq c+1$, $x<y$, and (3.3) holds. ∎
We now come to the most technical part of the proof, where given any $c\geq 500$ and $2c\leq b\leq\frac{c^{2}}{2}$, we will find a suitable integer $q$. We first make a remark about this inequality.
@@
To minimize this, we want $q$ to be large, and to have $x\geq c+1$, we could take $q\approx\frac{b}{c}$, which would give us $\left\lceil\frac{x-1}{y-x}\right\rceil x\approx b$, so this is promising. However, the floors and ceilings mean that the denominator $(y-x)$ may be closer to $(\frac{c}{q}-2)$, and this $(-2)$ term can really hurt us if $q$ is close to $\frac{c}{2}$. Therefore, when $b$ is close to $\frac{c^{2}}{2}$, we will need to make a clever choice of $q$.
@@
Fix positive integers $b$ and $c$ with $c\geq 500$ and $2c\leq b\leq\frac{c^{2}}{2}$. There exists a positive integer $q$ such that, letting $x$ and $y$ be as in (3.1), we have $x\geq c+1$, $x<y$, and the inequality (3.3) holds.
@@
$q=\frac{25}{26}(q+1)\geq\frac{25}{26}\frac{b}{c}.$
Because $b\leq\frac{c^{2}}{4.2}$, we have $q\leq\frac{b}{c}\leq\frac{c}{4.2}$. Now since $y-x$ is an integer and $y-x\geq\frac{c}{q}-2\geq 2.2$, we must have $y-x\geq 3$ and
@@
so because $y-x$ is an integer and $y-x\geq\frac{c}{q}-2>2.008$, we must have $y-x\geq 3$ and
@@
Also note that because $c\geq 500$, we have $\sqrt{b}\geq\frac{500}{\sqrt{3.4}}\geq 270$ and $q_{0}\geq\frac{270}{\sqrt{3.5}}-1\geq 140$. Let us first consider $q=q_{0}$. Use integer division to write $b+1=c_{0}q_{0}+r_{0}$ for some integers $c_{0}$ and $10\leq r_{0}\leq q_{0}-$ 1. If
@@
so at least three multiples of $(q_{0}-1)$, namely $(c_{0}+4)(q_{0}-1)$, $(c_{0}+5)(q_{0}-1)$, and $(c_{0}+6)(q_{0}-1)$, lie strictly between $(b+1)$ and $(b+c)$. This means that $x\leq c_{0}+4$ and $y\geq c_{0}+6$, so $y-x\geq 2$. This proves our claim that one of $q=q_{0}$ or $q=q_{0}-1$ will have $y-x\geq 2$.
@@
Now we can prove the remaining Part 3 of Theorem 3.1.
Proposition 3.18.
If $c\geq 2$ and $2c\leq b\leq\frac{c^{2}}{2}$, then $G$ is missing a connected partition of some type $\lambda\vdash n$.
Proof.
This follows directly from Proposition 3.13, Proposition 3.15, Lemma 3.14, and Lemma 3.17. ∎
@@
Because $S$ has four legs, we have $c\geq c_{1}+1\geq 2$. If $b\leq\frac{c^{2}}{2}$, the result follows from Parts 1, 2, and 3 of Theorem 3.1. If $b\geq\frac{c^{2}}{2}$, then
@@
and Zheng showed [spiderepos, Corollary 4.6] that such a spider $S$ is not $e$-positive. ∎
In order to prove Conjecture 1.1, it now remains to consider trees with a vertex of degree $4$ that is adjacent to a leaf, such as the spider $S(6,4,1,1)$. It turns out that $S(6,4,1,1)$ has a connected partition of every type $\lambda\vdash 13$, which shows that Wolfgang’s necessary condition will not be enough. We conclude by showing that there are infinitely many such examples.
Figure 3. The spider $S(6,4,1,1)$ has a connected partition of every type $\lambda\vdash 13$
$S(6,4,1,1)$
@@
Let $m\geq 1$ be an integer. The spider $S(6m,6m-2,1,1)$ has a connected partition of type $\lambda$ for every $\lambda\vdash 12m+1$.
@@
We claim that if $\lambda=(\lambda_{1},\ldots,\lambda_{\ell})$ has a rearrangement $\alpha$ with $6m-1,6m\notin\text{set}(\alpha)$, then $S(6m,6m-2,1,1)$ has a connected partition of type $\lambda$. Label the vertices on the leg of length $26m-$ 2 as $u_{1},\ldots,u_{6m-2}$, moving in from the leaf, and label the vertices on the leg of length $6m$ as $v_{1},\ldots,v_{6m}$ similarly. Suppose that $\alpha$ is a rearrangement of $\lambda$ that satisfies
$\alpha_{1}+\cdots+\alpha_{i}\leq 6m-2\text{ and }\alpha_{1}+\cdots+\alpha_{i+1}\geq 6m+1.$
Then we can partition the vertices $\{u_{1},u_{2},\ldots,u_{\alpha_{1}+\cdots+\alpha_{i}}\}$ into paths of sizes $\alpha_{1},\ldots,\alpha_{i}$, we can partition the vertices $\{v_{1},v_{2},\ldots,v_{\alpha_{i+2}+\cdots+\alpha_{\ell}}\}$ into paths of sizes $\alpha_{i+2},\ldots,\alpha_{\ell}$, and the remaining vertices form a connected set of size $\alpha_{i+1}$. This establishes our claim.
Note that if $\lambda$ has at least two $1$’s or if $\lambda=(2,2,\ldots,2,1)$, then $S(6m,6m-2,1,1)$ has a connected partition of type $\lambda$, so we may assume that $\lambda$ has at most one $1$ and has a part $x\geq 3$. By our claim, it suffices to find a rearrangement $\alpha$ of $\lambda$ with $6m-1,6m\notin\text{set}(\alpha)$. Fix a rearrangement $\beta$ of $\lambda$. We may assume that $16m-$ 1 or $6m\in\text{set}(\beta)$, otherwise we could take $\alpha=\beta$. Letting $\beta^{*}=(\beta_{\ell},\ldots,\beta_{1})$ denote the reverse composition, we may assume that $16m-$ 1 or $6m\in\text{set}(\beta^{*})$, otherwise we could take $\alpha=\beta^{*}$. Now there are only a few cases to consider.
@@
has $\text{set}(\alpha)=\{\ldots,6m-x-1,6m-x+1,6m+1,\ldots\}$, so $6m-1,6m\notin\text{set}(\alpha)$. Otherwise, the parts $(\beta_{1},\ldots,\beta_{i})$ consist of a single $1$ and $(3m-1)$ $2$’s, some $\beta_{k}=x\geq 3$ for $i+2\leq k\leq\ell$, and the composition
@@
has $\text{set}(\alpha)=\{\ldots,6m-1-y,6m+2-y,6m+2,\ldots\}$, so $6m-1,6m\notin\text{set}(\alpha)$. Now we may assume that every $1\leq k\leq i$ has $\beta_{k}=2$ or $3$, and by reversing, that every $i+2\leq k\leq\ell$ also has $\beta_{k}=2$ or $3$. Because $16m-$ 1 is not divisible by either $2$ or $3$, there must be some $1\leq k\leq i$ with $\beta_{k}=3$ and some $i+2\leq k^{\prime}\leq\ell$ with $\beta_{k^{\prime}}=2$. Now the composition

### v2

@@ 3. Proof of main result
By Lemma 3.10, there exists a partition $\lambda$ of $n$ with all parts in the interval $J=\{b+1,\ldots,b+c\}$. Now by Lemma 3.9 with $q=1$, $G$ is missing a connected partition of type $\lambda$. ∎
We can now prove Part (1) of Theorem 3.2, the case where $b\leq 2c-2$; and Part (2) of Theorem 3.2, the case where $c\geq c_{1}+1$ and b=2c1b=2c-1.
Proof of Part (1) of Theorem 3.2. .
Because $b\leq 2c-2$, we have
so the result follows from Lemma 3.12. ∎
Proof of Part (2) of Theorem 3.2. .
First note that if $c=2$, then $b=3$, and taking $\lambda=(2,\ldots,2)$ if $n$ is even or $\lambda=(3,2,\ldots,2)$ if $n$ is odd, every rearrangement $\alpha$ has the partial sum $\alpha_{1}+\alpha_{2}$ in the interval $I=\{4,5\}$ and we are done by Lemma 3.7. Now suppose that $c\geq 3$. Let $x=\left\lceil\frac{b+1}{2}\right\rceil=c\geq c_{1}+1$ and $y=\left\lfloor\frac{b+c}{2}\right\rfloor=\left\lfloor\frac{3c-1}{2}\right\rfloor$. By Lemma 3.10, we can find a partition $\lambda$ of $n$ with all parts in the interval $J=\{x,\ldots,y\}$ because
By Lemma 3.9 with $q=2$, $G$ is missing a connected partition of type $\lambda$. ∎
Our last case to prove is Part (3) of Theorem 3.2, the case where $2c\leq b\leq\frac{c^{2}}{2}$. Our strategy is as follows. Given integers $b$, $c$, and $n$ with $c\geq 2$, $2c\leq b\leq\frac{c^{2}}{2}$, and $n\geq 2b+c+1$, we wish to find a positive integer $q$ such that the integers
(3.3) $x=\left\lceil\frac{b+1}{q}\right\rceil\text{ and }y=\left\lfloor\frac{b+c}{q}\right\rfloor$
satisfy that $x\geq c+1\geq c_{1}+1$, $x<y$, and that there exists a partition $\lambda$ of $n$ with all parts in the interval $J=\{x,\ldots,y\}$. Then $G$ will have no connected partition of type $\lambda$ by Lemma 3.9. By Lemma 3.10, a sufficient condition to find such a $\lambda\vdash n$ is if
(3.4) $\left\lceil\frac{x-1}{y-x}\right\rceil x\leq 2b+c+1.$
Example 3.13.
Let $b=75$ and $c=14$, so that $n\geq 2b+c+1=165$.
  • $\circ$
    For $q\geq 6$, we have $x\leq\left\lceil\frac{76}{6}\right\rceil=13\leq c$, so this will not work.
  • $\circ$
    For $q=5$, we have $x=\left\lceil\frac{76}{5}\right\rceil=16$ and $y=\left\lfloor\frac{89}{5}\right\rfloor=17$. We have $x\geq c+1$, but
    $\left\lceil\frac{x-1}{y-x}\right\rceil x=240>165,$
    so this will not work. We cannot write $n=239$ as a sum of $16$’s and $17$’s.
  • $\circ$
    For $q=4$, we have $x=\left\lceil\frac{76}{4}\right\rceil=19$ and $y=\left\lfloor\frac{89}{4}\right\rfloor=22$. We have $x\geq c+1$ and
    $\left\lceil\frac{x-1}{y-x}\right\rceil x=114\leq 165,$
    so this will work. Lemma 3.10 tells us that there exists a partition $\lambda$ of $n$ with all parts in the interval $J=\{19,20,21,22\}$ and Lemma 3.9 tells us that $G$ will be missing a connected partition of type $\lambda$.
The next example shows that there may be no single value of $q$ for which the inequality (3.4) holds. However, for every value of $n\geq 2b+c+1$, there is some $q$ such that there exists $\lambda\vdash n$ with all parts in the interval $J=\{x,\ldots,y\}$, where $x$ and $y$ are as in (3.3).
Example 3.14.
Let $b=24$ and $c=7$, so that $n\geq 2b+c+1=56$.
  • $\circ$
    For $q\geq 4$, we have $x\leq 7\leq c$, so this will not work.
  • $\circ$
    For $q=3$, we have $x=9$, $y=10$, and $\left\lceil\frac{x-1}{y-x}\right\rceil x=72>56$, so this will not work.
  • $\circ$
    For $q=2$, we have $x=13$, $y=15$, and $\left\lceil\frac{x-1}{y-x}\right\rceil x=78>56$, so this will not work.
  • $\circ$
    For $q=1$, we have $x=25$, $y=31$, and $\left\lceil\frac{x-1}{y-x}\right\rceil x=100>56$, so this will not work.
We find that no single value of $q$ satisfies (3.4). For $n\geq 72$, we can take $q=3$, but this will not work for every $n\geq 56$. Nevertheless, for each $56\leq n\leq 71$, there is some $q$ such that there exists $\lambda\vdash n$ with all parts in the interval $J=\{x,\ldots,y\}$, where $x$ and $y$ are as in (3.3).
$q=3$ $:\{\underbrace{54,55,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}56},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}57},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}58},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}59},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}60}}_{6J}\}\cup\{\underbrace{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}63},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}64},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}65},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}66},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}67},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}68},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}69},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}70}}_{7J}\}\cup\{\underbrace{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}72},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}73},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}74},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\ldots}}_{8J}\}\cup\cdots$
$q=2$ $:\{\underbrace{52,53,54,55,56,57,58,59,60}_{4J}\}\cup\{\underbrace{65,66,67,68,69,70,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}71},72,73,74,75}_{5J}\}\cup\cdots$
$q=1$ $:\{\underbrace{50,51,52,53,54,55,56,57,58,59,60,{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}61},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}62}}_{2J}\}\cup\{\underbrace{75,76,77,\ldots}_{3J}\}\cup\cdots$
We will soon show by an analytic argument that if $c\geq 500$ and $2c\leq b\leq\frac{c^{2}}{2}$, then we can find a positive integer $q$ such that the integers $x$ and $y$ as in (3.3) satisfy $x\geq c+1$, $x<y$, and the inequality (3.4). Our estimates will not be valid for $2\leq c\leq 499$, so we check these cases by computer. For $41\leq c\leq 499$, we can always find a single value of $q$ for which (3.4) holds, as in Example 3.13. When $2\leq c\leq 40$, we will sometimes need different values of $q$ depending on $n$ to find a suitable partition $\lambda$ of $n$, as in Example 3.14. Sage code used to check the following Lemma will be provided in the Appendix.
Lemma 3.15.
Fix positive integers $b$, $c$, and $n$ with $2\leq c\leq 499$, $2c\leq b\leq\frac{c^{2}}{2}$, and $n\geq 2b+c+1$. There exists a positive integer $q$ such that the integers $x$ and $y$ as in (3.3) satisfy that $x\geq c+1$, $x<y$, and there exists a partition $\lambda$ of $n$ with all parts in the interval $J=\{x,\ldots,y\}$.
@@
First suppose that $2\leq c\leq 40$. By Lemma 3.10, we can take $q=1$ if $n\geq\left\lceil\frac{b}{c-1}\right\rceil(b+1)$, so there are finitely many triples $(b,c,n)$ to consider. For each one, we checked by computer, using the Sage code in Method 4.1, that there exists such a positive integer $q$.
Now suppose that $41\leq c\leq 499$. By Lemma 3.10, there exists such a partition $\lambda$ if the inequality (3.4) holds. There are finitely many pairs $(b,c)$ to consider. For each one, we checked by computer, using the Sage code in Method 4.2, that there exists such a positive integer $q$. ∎
We now come to the most technical part of the proof, where given any $c\geq 500$ and $2c\leq b\leq\frac{c^{2}}{2}$, we will find an integer $q$ such that the integers $x$ and $y$ as in (3.3) satisfy $x\geq c+1$, $x<y$, and the inequality
@@
We first make a remark about this inequality.
@@
To minimize this, we want $q$ to be large, and to have $x\geq c+1$, we could take $q\approx\frac{b}{c}$, which would give us $\left\lceil\frac{x-1}{y-x}\right\rceil x\approx b$, so this is promising. However, the floors and ceilings mean that the denominator (yx)(y-x) may be closer to $(\frac{c}{q}-2)$, and this (2)(-2) term can really hurt us if $q$ is close to $\frac{c}{2}$. Therefore, when $b$ is close to $\frac{c^{2}}{2}$, we will need to make a clever choice of $q$.
@@
Fix positive integers $b$ and $c$ with $c\geq 500$ and $2c\leq b\leq\frac{c^{2}}{2}$. There exists a positive integer $q$ such that, letting
(3.6) $x=\left\lceil\frac{b+1}{q}\right\rceil\text{ and }y=\left\lfloor\frac{b+c}{q}\right\rfloor,$
we have $x\geq c+1$, $x<y$, and the inequality
(3.7) $\left\lceil\frac{x-1}{y-x}\right\rceil x\leq 2b+c+1.$
@@
$q\geq\frac{25}{26}(q+1)\geq\frac{25}{26}\frac{b}{c}.$
Because $b\leq\frac{c^{2}}{4.2}$, we have $q\leq\frac{b}{c}\leq\frac{c}{4.2}$. Now since yxy-x is an integer and $y-x\geq\frac{c}{q}-2\geq 2.2$, we must have $y-x\geq 3$ and
@@
so because yxy-x is an integer and $y-x\geq\frac{c}{q}-2>2.008$, we must have $y-x\geq 3$ and
@@
Also note that because $c\geq 500$, we have $\sqrt{b}\geq\frac{500}{\sqrt{3.4}}\geq 270$ and $q_{0}\geq\frac{270}{\sqrt{3.5}}-1\geq 140$. Let us first consider $q=q_{0}$. Use integer division to write $b+1=c_{0}q_{0}+r_{0}$ for some integers $c_{0}$ and $0\leq r_{0}\leq q_{0}-1$. If
@@
so at least three multiples of $(q_{0}-1)$, namely $(c_{0}+4)(q_{0}-1)$, $(c_{0}+5)(q_{0}-1)$, and $(c_{0}+6)(q_{0}-1)$, lie strictly between $b+1$ and $b+c$. This means that $x\leq c_{0}+4$ and $y\geq c_{0}+6$, so $y-x\geq 2$. This proves our claim that one of $q=q_{0}$ or $q=q_{0}-1$ will have $y-x\geq 2$.
@@
Now we can prove the remaining Part (3) of Theorem 3.2, the case where $2c\leq b\leq\frac{c^{2}}{2}$.
Proof of Part (3) of Theorem 3.2. .
If $2\leq c\leq 499$, then Lemma 3.15 tells us that there is an integer $q$ and a partition $\lambda$ of $n$ with all parts in the corresponding interval $J=\{x,\ldots,y\}$, where $x$ and $y$ are as in (3.3). If $c\geq 500$, then Lemma 3.17 and Lemma 3.10 again give us these $q$ and $\lambda\vdash n$. Either way, now Lemma 3.9 tells us that $G$ is missing a connected partition of this type $\lambda\vdash n$. ∎
@@
Let $S$ be a spider with four legs, so that $c\geq c_{1}+1\geq 2$. If $b\leq\frac{c^{2}}{2}$, then the result follows from Parts (1), (2), and (3) of Theorem 3.2. If $b\geq\frac{c^{2}}{2}$, then
@@
and Zheng showed [19, Corollary 4.6] that such a spider $S$ is not $e$-positive. ∎
In order to prove Conjecture 1.2, it now remains to consider trees with a vertex of degree $4$ that is adjacent to a leaf, such as the spider $S(6,4,1,1)$ from Figure 1. It turns out that $S(6,4,1,1)$ has a connected partition of every type $\lambda\vdash 13$, which shows that Wolfgang’s necessary condition will not be enough. We conclude by showing that there are infinitely many such examples.
@@
Let $m\geq 1$ be an integer. The spider S(6m,6m2,1,1)S(6m,6m-2,1,1) has a connected partition of type $\lambda$ for every $\lambda\vdash 12m+1$.
Note that by Theorem 1.6, the spider S(6m,6m2,1,1)S(6m,6m-2,1,1) is not $e$-positive.
@@
We claim that if $\lambda=(\lambda_{1},\ldots,\lambda_{\ell})$ has a rearrangement $\alpha$ with $6m-1,6m\notin\text{set}(\alpha)$, then S(6m,6m2,1,1)S(6m,6m-2,1,1) has a connected partition of type $\lambda$. Label the vertices on the leg of length 6m26m-2 as $u_{1},\ldots,u_{6m-2}$, moving in from the leaf, and label the vertices on the leg of length $6m$ as $v_{1},\ldots,v_{6m}$ similarly. Suppose that $\alpha$ is a rearrangement of $\lambda$ that satisfies
[…]
Then we can partition the vertices $\{u_{1},u_{2},\ldots,u_{\alpha_{1}+\cdots+\alpha_{i}}\}$ into paths of orders $\alpha_{1},\ldots,\alpha_{i}$, we can partition the vertices $\{v_{1},v_{2},\ldots,v_{\alpha_{i+2}+\cdots+\alpha_{\ell}}\}$ into paths of orders $\alpha_{i+2},\ldots,\alpha_{\ell}$, and the remaining vertices form a connected set of order $\alpha_{i+1}$. This establishes our claim.
Note that if $\lambda$ has at least two $1$’s or if $\lambda=(2,2,\ldots,2,1)$, then S(6m,6m2,1,1)S(6m,6m-2,1,1) has a connected partition of type $\lambda$, so we may assume that $\lambda$ has at most one $1$ and has a part $x\geq 3$. By our claim, it suffices to find a rearrangement $\alpha$ of $\lambda$ with $6m-1,6m\notin\text{set}(\alpha)$. Fix a rearrangement $\beta$ of $\lambda$. We may assume that 6m16m-1 or $6m\in\text{set}(\beta)$, otherwise we could take $\alpha=\beta$. Letting $\beta^{*}=(\beta_{\ell},\ldots,\beta_{1})$ denote the reverse composition, we may assume that 6m16m-1 or $6m\in\text{set}(\beta^{*})$, otherwise we could take $\alpha=\beta^{*}$. Now there are only a few cases to consider.
@@
has $\text{set}(\alpha)=\{\ldots,6m-x-1,6m-x+1,6m+1,\ldots\}$, so $6m-1,6m\notin\text{set}(\alpha)$. Otherwise, the parts $(\beta_{1},\ldots,\beta_{i})$ consist of a single $1$ and (3m1)(3m-1) 22’s, some $\beta_{k}=x\geq 3$ for $i+2\leq k\leq\ell$, and the composition
@@
has $\text{set}(\alpha)=\{\ldots,6m-1-y,6m+2-y,6m+2,\ldots\}$, so $6m-1,6m\notin\text{set}(\alpha)$. Now we may assume that every $1\leq k\leq i$ has $\beta_{k}=2$ or $3$, and by reversing, that every $i+2\leq k\leq\ell$ also has $\beta_{k}=2$ or $3$. Because 6m16m-1 is not divisible by either $2$ or $3$, there must be some $1\leq k\leq i$ with $\beta_{k}=3$ and some $i+2\leq k^{\prime}\leq\ell$ with $\beta_{k^{\prime}}=2$. Now the composition
@@
4. Appendix
Method 4.1.
This is used to check that for integers $b$, $c$, and $n$ with $2\leq c\leq 40$, $2c\leq b\leq\frac{c^{2}}{2}$, and $2b+c+1\leq n\leq\left\lceil\frac{b}{c-1}\right\rceil(b+1)$, we can find a positive integer $q$ such that the integers $x$ and $y$ as in (3.3) satisfy $x\geq c+1$, $x<y$, and $tx\leq n\leq ty$ for some $t$; and therefore that there exists a partition $\lambda$ of $n$ with all parts in the interval $J=\{x,\ldots,y\}$.
1def checkc40():
2 for c in range(2,41):
3 for b in range(2*c,floor(c^2/2)+1):
4 for n in range(2*b+c+1,ceil(b/(c-1))*(b+1)+1):
5 print(’b:’,b,’c:’,c,’n:’,n)
6 foundq=false
7 for q in range(1,ceil((b+1)/c)):
8 x=ceil((b+1)/q)
9 y=floor((b+c)/q)
10 t=floor(n/x)
11 if x>=c+1 and x<y and t*x<=n<=t*y:
12 print(’q:’,q,’x:’,x,’y:’,y,’t:’,t)
13 foundq=true
14 if not foundq:
15 return false
16 return true
Method 4.2.
This is used to check that for integers $b$ and $c$ with $2\leq c\leq 40$ and $2c\leq b\leq\frac{c^{2}}{2}$, we can find a positive integer $q$ such that the integers $x$ and $y$ as in (3.3) satisfy $x\geq c+1$, $x<y$, and the inequality (3.4); and therefore that there exists a partition $\lambda$ of $n$ with all parts in the interval $J=\{x,\ldots,y\}$.
1def checkc500():
2 for c in range(41,500):
3 for b in range(2*c,floor(c^2/2)+1):
4 print(’b:’,b,’c:’,c)
5 foundq=false
6 for q in range(1,ceil((b+1)/c)):
7 x=ceil((b+1)/q)
8 y=floor((b+c)/q)
9 M=ceil((x-1)/(y-x))*x
10 if x>=c+1 and x<y and M<=2*b+c+1:
11 print(’q:’,q,’x:’,x,’y:’,y,’M:’,M)
12 foundq=true
13 if not foundq:
14 return false
15 return true